OpenStudy (anonymous):

How is 8x/(x-2) simplified to 8/(1-(2/x))? This is from PS1 1D-1 letter g.

OpenStudy (fifciol):

\[\frac{ 8x }{ x-2 }=\frac{ 8x }{x(1-\frac{ 2 }{ x} )}=\frac{ 8 }{1-\frac{ 2 }{ x } }\]

OpenStudy (anonymous):

Thanks! I had figured it out already, but at least your reply confirms what I got. I should have just played around a little bit with the algebra before posting :)